i want to ask how to solve this problems . i want u to explain me the whole question n then give a solution to it. my ques is
show that my=x^2-4(x-1) meets the curve y=x^2-3x+2 at two distinct points for all non zero values of m.i would be very grateful to u if u solve this ques for me.this ques is nt from text book.






Answer by stanbon(75887)

show that my=x^2-4(x-1) meets the curve y=x^2-3x+2 at two distinct points
---------------------
y = x^2-4x+4
y = x^2-3x+2
------------------
Substitute to get:
x^2-4x+4 = x^2-3x+2
x= 2
----------
Substitute to solve for y:
y = 2^2-4*2+4
y = 0
----------
The curves meet at the point (2,0)
-----------
Cheers,
Stan H.